# Permutations, substitutions and finite axiomatizability

### v1

@@ 1. What was and what is to come
$\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x,\quad\mathsf{s}_{ii}(x)=x,$ (7)
@@
$\mathsf{s}_{ij}x=\mathsf{c}_{i}\mathsf{s}_{ij}x.$ (9)
@@
$\mathsf{p}_{ij}\mathsf{p}_{ij}x=x,\quad\mathsf{p}_{ii}x=x,\quad\mathsf{p}_{ij}x=\mathsf{p}_{ji}x,$ (11)
$\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x,$ (12)
$\mathsf{p}_{ij}\mathsf{s}_{ij}x=\mathsf{s}_{ji}\mathsf{p}_{ij}x.$ (13)
@@ 4. Finitary polyadic algebras and substitution-cylindric algebras
The identities 8-(F9) are positive equations. It is only (F6) which needs a comment. The equation $\mathsf{s}_{ij}(x)=\mathsf{s}_{ij}(-x)$ is part of what it means that $\mathsf{s}_{ij}$ is a Boolean endomorphism, and it is visible that the negation symbol is used here. This equation, as noted in [22, p.561], can be replaced by the ones expressing that $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$. This is because if $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$, then
@@
In Section 6, we will use the following consequences of the $\mathsf{FPA}_{\alpha}$-axioms.
@@
(P4) and (P5) are proved as Lemma 1.6 in [22]. [^9] We note that there are two typos in the statement of Lemma 1.6 in [22]. Namely, (i) writes $\mathsf{p}_{ij}\mathsf{c}_{i}=\mathsf{c}_{i}\mathsf{p}_{ij}$, but the correct version is our (P4). In (iii), the correct assumption is that $\sigma(i)=i$ and for $j\neq i$, $\sigma(j)\neq i$. However, the proof of (the typo-free version of) Lemma 1.6 in [22] is correct, and throughout that paper the lemma is used with the correct statement. $\blacksquare$
Recall that permutations $\sigma:\alpha\to\alpha$ are generated by the transpositions. We let $\mathsf{p}_{[i,j]}=\mathsf{p}_{ij}$ and $\mathsf{p}_{\sigma\circ\tau}=\mathsf{p}_{\sigma}\circ\mathsf{p}_{\tau}$. This way to every permutation $\sigma$ we assign a corresponding polyadic term $\mathsf{p}_{\sigma}$. [22, Claim 1.2] shows that different decompositions of the same permutation into products of transpositions yield identical terms (in $\mathsf{FPA}_{\alpha}$). For a precise definition we refer to [22, p.554]. This leads to the important lemma below that we make use of many times in Section 6.
@@
Next we turn to algebras without the $\mathsf{p}_{ij}$ operations. The class $\mathsf{SCA}_{\alpha}$ of algebras is axiomatized by the finitary polyadic algebra axioms of $\mathsf{FPA}_{\alpha}$ without the $\mathsf{p}_{ij}$’s, that is, when the $\mathsf{FPA}_{\alpha}$ axioms speaking about $\mathsf{p}_{ij}$ are deleted. We will see that these algebras are obtained from cylindric algebras by replacing the diagonal constants $\mathsf{d}_{ij}$ with the derived operations $\mathsf{s}_{ij}$. The elements of $\mathsf{SCA}_{\alpha}$ will be called substitution-cylindric algebras.
Definition 4.5 (Cf. Sax on p.559 in [22]).
@@
Proof . The proof of (i) relies on results in Pinter [19] and Sain-Thompson [22]. First we prove that (F0)-(F6) imply substitution axioms (S1)-(S6) in [19]. Indeed, Lemma 1.5(iii) and Proposition 6 in [22] prove that the equations
  1. (A1)
    $\mathsf{s}_{ij}\mathsf{s}_{ki}x=\mathsf{s}_{ij}\mathsf{s}_{kj}x$
@@
hold in $\mathsf{SCA}_{\alpha}$ and (S1)-(S6) are (F6), (F0), (A1), (A2), and (A3). Also, [22] proves
@@
To prove (ii), notice first that $\mathsf{SCA}_{\alpha}\subseteq\mathbf{S}\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$ follows from (i) together with $\mathsf{SCA}_{\alpha}$ being a canonical variety. For the other direction, we have to prove that (F0)-(F6) hold in $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$. Now, (F0) is (C0)+[12, 1.5.1], (F1) is (C2), and (F2)-(F6) are [12, 1.2.6(ii), 1.5.8(i), 1.5.9(ii), 1.5.8(ii), 1.5.3], respectively. $\blacksquare$
@@ 5. $\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}^{c}_{\alpha}$
$\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$ (by the equations (F0)-(F6) defining $\mathsf{SCA}_{\alpha}$).
@@
We now prove Theorem 1.2. Recall first that $\mathsf{Rdf}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and $\mathsf{RFPA}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$. Then $\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}=\mathbf{S}\mathbf{SP}\operatorname{Rd}_{cs}\mathsf{RA}_{\alpha}^{csp}=\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$. So Theorem 1.2 states that $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ iff ($\operatorname{Rd}_{c}\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and equations (6)–(9) hold in $\mathfrak{A}$). By Theorem 5.10 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{c}$ and equations (F0)-(F6) hold in $\mathfrak{A}$). Equations (6)–(9) in Theorem 1.2 are the equations in (F0)-(F6) in which the substitutions occur [^11] The indices $i,j,k$ in Theorem 1.2 are required to be distinct while in (F0)–(F6) they are not required to be distinct. It is easy to check that this does not make any difference.; the ones in which the substitutions do not occur hold in $\mathsf{RA}_{\alpha}^{c}$, so they are superfluous. Theorem 1.2 has been proved.
@@ 6. $\mathsf{RA}_{\alpha}^{csp}$ is finitely axiomatizable over $\mathsf{RA}^{cs}_{\alpha}$
Proof of Lemma 6.2. For any $i,j<\alpha$, from (83), [12, 1.5.4(i)] and cylindric axiom (C2) we get $\mathop{S}\nolimits_{ij}^{U}(\mathsf{d}_{ij}^{\mathfrak{A}})=1$ and $x\cap\mathsf{d}_{ij}^{\mathfrak{A}}\subseteq\mathop{S}\nolimits_{ij}^{U}(x)$. Sain–Thompson [22] define the class $\mathsf{FPEA}_{\alpha}$ as $\mathsf{FPA}_{\alpha}$ enriched with diagonal elements $\mathsf{d}_{ij}$ satisfying $\mathsf{s}_{ij}\mathsf{d}_{ij}=1$ and $x\cdot\mathsf{d}_{ij}\leq\mathsf{s}_{ij}x$ (see [22, p.544, Definition 2]). We just verified that these two properties hold for $\mathop{S}\nolimits_{ij}^{U}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, therefore the expanded structure $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is an $\mathsf{FPEA}_{\alpha}$. [22] proves that $\mathsf{FPEA}_{\alpha}$ are term equivalent to quasi-polyadic equality algebras $\mathsf{QPEA}_{\alpha}$, and in particular, [22, p.557-558, (E3)] proves that from $\mathsf{FPEA}_{\alpha}$ the identity
@@
We prove Theorem 1.1. By Theorem 1.2 and Theorem 1.3 we get that $\mathsf{RFPA}_{\alpha}$ is axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations (F0)–(F9) defining $\mathsf{FPA}_{\alpha}$. [22, Theorem 1(i)] states that $\mathsf{FPA}_{\alpha}$ is term-definitionally equivalent to $\mathsf{PA}_{\alpha}$. The same term-definitions take $\mathsf{RFPA}_{\alpha}$ to $\mathsf{RPA}_{\alpha}$. Hence, $\mathsf{RPA}_{\alpha}$ is finitely axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations defining $\mathsf{PA}_{\alpha}$. Examining the defining equations of $\mathsf{PA}$, in e.g., [12, Def.5.4.1], one sees that among these equations (1)–(4) are the ones in which the substitution operations $\mathsf{s}_{\sigma}$ occur and the rest follow easily from (F0)–(F9). Theorem 1.1 has been proved.

### v2

@@ 1. What was and what is to come
$\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x,\quad\mathsf{s}_{ij}x=\mathsf{c}_{i}\mathsf{s}_{ij}x,\quad\mathsf{s}_{ii}(x)=x,$ (7)
@@
$\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x.$ (9)
@@
$\mathsf{p}_{ij}\mathsf{p}_{ij}x=x,\quad\mathsf{p}_{ij}x=\mathsf{p}_{ji}x,\quad\mathsf{p}_{ii}x=x,$ (11)
$\mathsf{p}_{ij}\mathsf{s}_{ij}x=\mathsf{s}_{ji}\mathsf{p}_{ij}x,$ (12)
$\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x.$ (13)
@@ 4. Finitary polyadic algebras and substitution-cylindric algebras
The identities 8-(F9) are positive equations. It is only (F6) which needs a comment. The equation $\mathsf{s}_{ij}(x)=\mathsf{s}_{ij}(-x)$ is part of what it means that $\mathsf{s}_{ij}$ is a Boolean endomorphism, and it is visible that the negation symbol is used here. This equation, as noted in [23, p.561], can be replaced by the ones expressing that $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$. This is because if $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$, then
@@
In Section 6, we will use consequences (P1)(P5) of the $\mathsf{FPA}_{\alpha}$-axioms. In Definition 4.5 below and in section 5 we will use consequence (FS) of the $\mathsf{FPA}_{\alpha}$-axioms that concerns solely the substitution operations $\mathsf{s}_{ij},i,j<\alpha$. It is proved in [23, Remark 7] that (FS) does not follow from the axioms of $\mathsf{FPA}_{\alpha}$ that do not contain any reference to $\mathsf{p}_{ij},i,j<\alpha$.
@@
  6. (FS)
    $\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x$ if $i,j,k<\alpha$ are distinct.
@@
(P4) and (P5) are proved as Lemma 1.6 in [23]. [^9] We note that there are two typos in the statement of Lemma 1.6 in [23]. Namely, (i) writes $\mathsf{p}_{ij}\mathsf{c}_{i}=\mathsf{c}_{i}\mathsf{p}_{ij}$, but the correct version is our (P4). In (iii), the correct assumption is that $\sigma(i)=i$ and for $j\neq i$, $\sigma(j)\neq i$. However, the proof of (the typo-free version of) Lemma 1.6 in [23] is correct, and throughout that paper the lemma is used with the correct statement. (FS) is proved as Lemma 1.5(iii) in [23]. $\blacksquare$
Recall that permutations $\sigma:\alpha\to\alpha$ are generated by the transpositions. We let $\mathsf{p}_{[i,j]}=\mathsf{p}_{ij}$ and $\mathsf{p}_{\sigma\circ\tau}=\mathsf{p}_{\sigma}\circ\mathsf{p}_{\tau}$. This way to every permutation $\sigma$ we assign a corresponding polyadic term $\mathsf{p}_{\sigma}$. [23, Claim 1.2] shows that different decompositions of the same permutation into products of transpositions yield identical terms (in $\mathsf{FPA}_{\alpha}$). For a precise definition we refer to [23, p.554]. This leads to the important lemma below that we make use of many times in Section 6.
@@
Next we turn to algebras without the $\mathsf{p}_{ij}$ operations. The class $\mathsf{SCA}_{\alpha}$ of algebras is axiomatized by the finitary polyadic algebra axioms of $\mathsf{FPA}_{\alpha}$ without the $\mathsf{p}_{ij}$’s, that is, when the $\mathsf{FPA}_{\alpha}$ axioms speaking about $\mathsf{p}_{ij}$ are deleted; with (FS) added as axiom. We will see that these algebras are obtained from cylindric algebras by replacing the diagonal constants $\mathsf{d}_{ij}$ with the derived operations $\mathsf{s}_{ij}$. The elements of $\mathsf{SCA}_{\alpha}$ will be called substitution-cylindric algebras. This class was introduced by Pinter [20] where it is denoted by $\mathsf{QA}_{\alpha}$, and its members are called quantifier-algebras.
Definition 4.5 (Cf. Definition 2.1 in [20]).
@@
  8. (FS)
    $\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x$ if $i,j,k<\alpha$ are distinct.
@@
Proof . The proof of (i) relies on results in Pinter [19] and Sain-Thompson [23]. First we prove that (F0)-(F6),(FS) imply substitution axioms (S1)-(S6) in [19]. Indeed, Proposition 6 in [23] proves that the equations
@@
hold in $\mathsf{SCA}_{\alpha}$ and (S1)-(S6) are (F6), (F0), (FS),(A1), and (A2). Also, [23] proves
@@
To prove (ii), notice first that $\mathsf{SCA}_{\alpha}\subseteq\mathbf{S}\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$ follows from (i) together with $\mathsf{SCA}_{\alpha}$ being a canonical variety. For the other direction, we have to prove that (F0)-(F6),(FS) hold in $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$. Now, (F0) is (C0)+[12, 1.5.1], (F1) is (C2), and (F2)-(F6),(FS) are [12, 1.2.6(ii), 1.5.8(i), 1.5.9(ii), 1.5.8(ii), 1.5.3, 1.5.10(ii)], respectively. $\blacksquare$
@@ 5. $\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}^{c}_{\alpha}$
$\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$ (by the equations (F0)-(F6),(FS) defining $\mathsf{SCA}_{\alpha}$).
@@
We now prove Theorem 1.2. Recall first that $\mathsf{Rdf}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and $\mathsf{RFPA}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$. Then $\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}=\mathbf{S}\mathbf{SP}\operatorname{Rd}_{cs}\mathsf{RA}_{\alpha}^{csp}=\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$. So Theorem 1.2 states that $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ iff ($\operatorname{Rd}_{c}\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and equations (6)–(9) hold in $\mathfrak{A}$). By Theorem 5.10 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{c}$ and equations (F0)-(F6),(FS) hold in $\mathfrak{A}$). Equations (6)–(9) in Theorem 1.2 are the equations in (F0)-(F6),(FS) in which the substitutions occur [^11] The indices $i,j,k$ in Theorem 1.2 are required to be distinct while in (F0)–(F6),(FS) they are not required to be distinct. It is easy to check that this does not make any difference.; the ones in which the substitutions do not occur hold in $\mathsf{RA}_{\alpha}^{c}$, so they are superfluous. Theorem 1.2 has been proved.
@@ 6. $\mathsf{RA}_{\alpha}^{csp}$ is finitely axiomatizable over $\mathsf{RA}^{cs}_{\alpha}$
Proof of Lemma 6.2. For any $i,j<\alpha$, from (83), [12, 1.5.4(i)] and cylindric axiom (C2) we get $\mathop{S}\nolimits_{ij}^{U}(\mathsf{d}_{ij}^{\mathfrak{A}})=1$ and $x\cap\mathsf{d}_{ij}^{\mathfrak{A}}\subseteq\mathop{S}\nolimits_{ij}^{U}(x)$. Sain–Thompson [23] define the class $\mathsf{FPEA}_{\alpha}$ as $\mathsf{FPA}_{\alpha}$ enriched with diagonal elements $\mathsf{d}_{ij}$ satisfying $\mathsf{s}_{ij}\mathsf{d}_{ij}=1$ and $x\cdot\mathsf{d}_{ij}\leq\mathsf{s}_{ij}x$ (see [23, p.544, Definition 2]). We just verified that these two properties hold for $\mathop{S}\nolimits_{ij}^{U}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, therefore the expanded structure $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is an $\mathsf{FPEA}_{\alpha}$. [23] proves that $\mathsf{FPEA}_{\alpha}$ are term equivalent to quasi-polyadic equality algebras $\mathsf{QPEA}_{\alpha}$, and in particular, [23, p.557-558, (E3)] proves that from $\mathsf{FPEA}_{\alpha}$ the identity
@@
We prove Theorem 1.1. By Theorems 1.21.3 and the last item (FS) in Lemma 4.3 we get that $\mathsf{RFPA}_{\alpha}$ is axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations (F0)–(F9) defining $\mathsf{FPA}_{\alpha}$. [23, Theorem 1(i)] states that $\mathsf{FPA}_{\alpha}$ is term-definitionally equivalent to $\mathsf{PA}_{\alpha}$. The same term-definitions take $\mathsf{RFPA}_{\alpha}$ to $\mathsf{RPA}_{\alpha}$. Hence, $\mathsf{RPA}_{\alpha}$ is finitely axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations defining $\mathsf{PA}_{\alpha}$. Examining the defining equations of $\mathsf{PA}$, in e.g., [12, Def.5.4.1], one sees that among these equations (1)–(4) are the ones in which the substitution operations $\mathsf{s}_{\sigma}$ occur and the rest follow easily from (F0)–(F9). Theorem 1.1 has been proved.